OpenStudy (anonymous):

90x^9y^2*3z^3 -------------- 15x^2y^7 show work please

OpenStudy (anonymous):

(18 x^7 z^3)/y^5

OpenStudy (anonymous):

i got this far 270x^9y^2z^3 -------------- 15x^2y^7

OpenStudy (anonymous):

how did you get that

OpenStudy (saifoo.khan):

simple cancellation

OpenStudy (anonymous):

well.. it is a multiplication. You can cut the numbers

OpenStudy (anonymous):

idk the english word for that but i believe it is what saifoo said.

OpenStudy (saifoo.khan):

hmmm.

OpenStudy (anonymous):

i guess if i knew what you were talking about I wouldn't have asked for help

OpenStudy (saifoo.khan):

LOL. u can say, its just multiplying and dividing by same number.. like in 6/4 the answer u get is 3/2 so u multiply & divide by 2

OpenStudy (anonymous):

90x^9y^2*3z^3 90x^9*3z^3 y^2 -------------- = ------------- x ------ 15x^2y^7 15x^2 y^7 isolating the second term: y^2 1 ------ = y^2 * y^(-7) = y^(2-7) = y^(-5) = ---- y^7 y^5 Now you just need to do this for "x" and "z". Understand now?

OpenStudy (anonymous):

LOL isolate the equation like this: 90x^9*3z^3 y^2 ----------- x ---------- 15x^2 y^7 if you take only on the second term: y^2 1 -------- = y^2 * y^(-7) = y^(2-7) = y^(-5) = ----- y^7 y^5 now you just need to do the same for the "x"